OpenStudy (stormswan):

Please help. I will give a medal and fan. I need help with these 5 questions. 1) m=5p+8 m=-10p+3 2) h=6g-4 h=-2g+28 3) y=x-2 2x+2y=4 4) 3x-6y=30 y=-6x+34 5) 7x-8y=112 y=-2x+9

OpenStudy (anonymous):

You can subtract the two equations to solve 1 and 2, and use substitution for the others.

OpenStudy (anonymous):

well i dont noe i am very weak in maths

OpenStudy (stormswan):

it says on my paper: Solve each system using substitution. Check your Solution.

OpenStudy (stormswan):

and I dont know how to do any of this, because I was sick and my teacher just gave me the homework. can someone please teach me how to do these?

OpenStudy (anonymous):

If you have to use substitution it would be easier to start with #3 then.

OpenStudy (anonymous):

You know that \(y=x-2\), so you can substitute that into the second equation to get \(2x + 2(x-2)=4\).

OpenStudy (stormswan):

ok but where does the y go then?

OpenStudy (anonymous):

\(y\) was replaced in the substitution because you have to get rid of it so that you can solve for \(x\). Once you find \(x\) you can plug that value back into one of the equations to find \(y\).

OpenStudy (stormswan):

oh, ok. so what do I do now that I have 2x+2(x-2)=4

OpenStudy (anonymous):

Solve for \(x\).

OpenStudy (stormswan):

so i would use pemdas? right

OpenStudy (anonymous):

You always need to respect order of operations but in this case you'll want to distribute the 2 in \(2(x-2)\) first.

OpenStudy (anonymous):

\[2x + 2(x-2)=4\]\[2x + 2x - 4 = 4\]\[4x=8\]\[x=2\]

OpenStudy (stormswan):

that sort of makes sense. how would I solve for y? @Richard_Feynman

OpenStudy (anonymous):

Plug the value of \(x\) you found back into the original equation with \(y\) and solve for \(y\).

OpenStudy (stormswan):

I dont understand... sorry.

OpenStudy (anonymous):

Look back at \(y=x+2\). You know that \(x = 2\), so \(y = 2 + 2\).

OpenStudy (stormswan):

ok, but how would I solve that?

OpenStudy (anonymous):

How would you solve what? \(y= 2 + 2\) trivially means that \(y=4\).

OpenStudy (anonymous):

There's nothing else to solve in #3 now that you have \(x\) and \(y\).

OpenStudy (stormswan):

but I meant how would I show my work and how would I actually set it up and solve it.